# Secure and Efficient Migration of Large Enclaves in a Data Center

### v4

@@ Abstract.
Cloud service providers are increasingly adopting Trusted Execution Environments, or TEEs, to provide hardware-guaranteed security to an application executing on remote, untrusted data centers. Often, there is a need to live-migrate such secure applications for load balancing or data center maintenance. Today, state-of-the-art migration methods for TEE still use the decade-old stop-and-copy-based method, which introduces large downtimes. This is because state-of-the-art live-migration approaches do not work for applications that run on TEEs.
We propose ConstMig, a near-zero downtime live-migration mechanism for large memory footprint TEE-based applications. We provide two alternatives: a kernel-based approach and a compiler-based approach. Based on the memory usage, we can prefer one approach over the other. ConstMig is fully compatible with containers, virtual machines (VMs) and microVMs. Our prototype, built on Intel SGX, a TEE solution from Intel, has a near-zero downtime irrespective of enclave size. ConstMig reduces the total downtime by 77-96% for a suite of SGX applications with multi-GB memory footprints compared to state-of-the-art TEE-based migration, MigSGX.
@@ 1. Introduction
Cloud service providers, or CSPs, are increasingly adopting trusted execution environments, or TEEs, to execute security-sensitive applications (Bacon et al., 2010; Orobosade et al., 2020; Singh and Chatterjee, 2017; Baumann et al., 2014; Gribov et al., 2017; Wang et al., 2020; Priebe et al., 2018; Antonopoulos et al., 2020; Dahl et al., 2018; Hong et al., 2020; tf encrypted, 2025; Tramer and Boneh, 2019; Shafran et al., 2021) on remote, untrusted cloud nodes. Leading CSPs such as Microsoft and Google provide several options for executing applications on such TEEs (Cfir Cohen, 2023; ju shim, 2023). Examples of such TEEs are Intel SGX (Intel, 2019c), Intel TDX (Cheng et al., 2024a), AMD SEV (Kaplan et al., 2016), ARM Trustone (Alves and Felton, 2004) and RISC-V Keystone (Lee et al., 2020). In a TEE setting, the hardware ensures the application’s security. The hypervisor and operating system (OS) are assumed to be untrusted (§2.2).
@@
However, managing a data center with TEEs presents unique challenges. In a data center, migration of an application or a virtual machine from one machine to another is a common and often a necessary operation (Khanna et al., 2006; Arzuaga and Kaeli, 2010; Zaw and Thein, 2012; Jin et al., 2011; Fu, 2009; Nagarajan et al., 2007; Troger and Salfner, 2011; Voorsluys et al., 2009; Stage and Setzer, 2009; Beloglazov and Buyya, 2010). CSPs rely on efficient live migration to provide high availability, load balancing, efficient resource management, failure tolerance, seamless upgradation, and code patching. In a non-TEE setting, modern migration mechanisms utilize efficient page access tracking mechanisms provided by modern hardware and operating systems to achieve a near-zero downtime during migration (Clark et al., 2005; Hines et al., 2009). There are two primary ways to achieve such live migration: pre-copy (Clark et al., 2005) and post-copy (Hines et al., 2009) migration methods (§2.1).
We investigated the problem of migrating secure applications on Intel’s modern hardware in this paper. We focused on Intel SGX (Intel, 2019c)2.2) which creates secure environments at the application level. Due to the security guarantees provided by SGX hardware, the conventional support required from hardware and OS, such as userspace fault handling and page write tracking, is not available (§3.3). Consequently, traditional live migration mechanisms cannot be directly applied to applications executing within SGX. Hence, prior work in enabling migration for SGX relied on the classical stop-and-copy method for migrating SGX enclaves (Nakashima and Kourai, 2021; Liang et al., 2019; Alder et al., 2018; Guerreiro et al., 2020; Soriente et al., 2019). In the stop-and-copy method, an executing application is paused, a checkpoint is taken, and the checkpoint is transferred to the destination machine, where it is subsequently resumed. This approach has a large downtime, which is proportional to the memory footprint (see Figure 1).
Confidential Virtual Machines: Intel Trusted Domain eXtension (TDX) and AMD Secure Encrypted Virtualization (SEV), the two latest TEE solutions from Intel and AMD (resp.), aim to provide confidential virtual machines (CVMs) (Misono et al., 2024; Kuvaiskii et al., 2024) to securely execute applications without modification. They both have added hardware support to enable live migration (Wang et al., 2021; AMD, 2022) (Nakashima and Kourai, 2021). Even though they appear to be the default solutions in this case, they provide security at the VM level (susceptible to attacks by co-hosted applications), they have a large TCB (susceptible to attacks), a high initialization time (increases downtime during migration), and a high memory encryption cost even in the unsecure portions of an application (performance overheads). SGX has a small TCB, low initialization time, and no overheads for the unsecure portion (§3.4 and Appendix B).
We present ConstMig a novel live migration mechanism for SGX enclaves with a near-zero downtime that is independent of the memory footprint of the application. ConstMig requires no additional hardware and has far lower performance overheads than the state-of-the-art (see Figure 1). It follows a three-step approach: ❶ we synchronously send a small part of the application’s runtime state to the destination machine at the outset. ❷ The application is resumed on the destination machine; ❸ the remaining memory pages are sent to the destination machine proactively (preference given to faulting pages) (§4).
@@
ConstMig addresses a fundamental challenge in enabling live migration for enclave-based systems: a fine-grained tracking of memory accesses on the destination machine to ensure that the restored application reads valid data from the source machine – instead of an invalid, all-zero page (§5). We propose two access tracking mechanisms: ❶ based on fine-grained page fault tracking ( §5.1) using kernel modifications and ❷ based on fine-grained pointer access tracking (§5.2). In addition to providing all the security guarantees of SGX, ConstMig also provides a freshness guarantee during the migration process, which ensures that an attacker cannot replay old encrypted enclave data during migration (§6). Our approach has been tested on real cloud settings and even works in container, VM, and microVM-based settings (Appendix D).
Contributions: Our key contributions are as follows:
@@
    We thoroughly characterize SGX and identify the key challenges faced while adopting modern migration schemes for enclaves (with minimal downtime).
@@
2. Background
2.1. Migration Methods
@@
The performance of a migration method is measured using two metrics: total downtime ($DT$) – total time the application is not running) and total migration time ($MT$) – duration from the start of the migration until its completion (Stoyanov and Kollingbaum, 2018; Nakashima and Kourai, 2021).
(a) Pre-copy migration
(b) Post-copy migration
@@
2.1.1. Stop-and-Copy Migration
In a stop-and-copy migration approach (Clark et al., 2005), the application being migrated is “paused” on the source machine, the state (registers, stack, heap, open files, network, etc.) is transferred to the destination machine, and then the application is resumed on the destination machine. The application remains paused during the state transfer, and hence, the total migration time and total downtime are proportional to the application’s memory footprint.
@@ 2.1.2. Pre-copy Migration
As shown in Figure  2a, the steps in pre-copy migration are as follows. ❶ “Snapshot” the state of the application and track subsequent write operations. ❷ Transfer the snapshot to the destination machine. ❸ Track the modified data pages (dirty pages). If the number of dirty pages is less than a threshold, pause the application. ❹ Transfer the dirty pages to the destination machine, and ❺ resume the application on the destination machine. However, if the number of dirty pages exceeds a threshold, only the dirty pages are transferred, and the process resumes from step ❸. Here, the total migration time is proportional to the memory footprint, and the total downtime is proportional to the number of dirty pages in the final tranche of pages that is sent.
@@ 2.1.3. Post-Copy Migration
In post-copy migration (see Figure 2b), ❶ the application is paused on the source machine, ❷ and a minimal amount of the state (CPU state, registers, and non-pageable memory (Hines et al., 2009; Clark et al., 2005)) is sent to the destination machine. Subsequently, ❸ the application is resumed on the destination machine. ❹ Next, we start the transfer of the application’s state from the source to the destination machine. During the state transfer, the migration method tracks the memory usage of the application on the destination machine. Any request to use data that is still on the source machine results in a network page fault, and is handled transparently by the migration mechanism and the OS. Here, the total migration time is proportional to the memory footprint, but the total downtime is constant. However, the application might slow down during the migration due to expensive network faults.
2.2. Intel Software Guard eXtensions or SGX
Intel Software Guard eXtension, or SGX, provides a secure way to execute an application on a remote, untrusted machine by creating a secure sandbox called an enclave. These enclaves are free from tampering even from privileged entities such as hypervisors, OS, or system administrators (Costan and Devadas, 2016). At boot time, SGX reserves a part of the main memory called the processor reserved memory (PRM). The PRM is managed by the hardware and is encrypted using a hardware key that is inaccessible to any entity, including the OS, hypervisor, SGX driver, and the SGX application. The application-usable portion of the PRM is called the enclave page cache (EPC) (the rest is used by SGX to store metadata).
In the SGX framework, an application is split into secure and unsecure portions. The secure portion executes within the enclave with associated performance overheads (see Figure 4a). The unsecure portion executes outside SGX in a conventional manner without any performance overhead. The unsecure portion can call a function inside an enclave using a mechanism called an enclave call or ECALL. Similarly, an enclave can also call functions in the unsecure component using a mechanism called outside call, or OCALL (Costan and Devadas, 2016).
Scalable SGX: The latest version of SGX, also known as scalable SGX, can support an EPC up to 512 GB (Intel, 2022b, a) compared to the erstwhile version, SGX (referred to as SGXv1 in the paper) which had a usable EPC of just 92 MB (Kumar et al., 2022; Kumar and Sarangi, 2021; Intel, 2019a). SGX guarantees that the enclave being executed has not been tampered with by validating the signature of a loaded enclave (Costan and Devadas, 2016) at runtime.
The initialization time of a large memory footprint enclave can be huge, as it requires all the pages to be added to it upfront. As per our measurements, initializing an enclave of 20 GB takes around $\approx$ 30 seconds. To solve this, Intel introduced Enclave Dynamic Memory Management, or EDMM (McKeen et al., 2016), which allows the user to configure the amount of memory that must be statically added during initialization – committed pages (Kumar et al., 2025). Additional data pages can be dynamically added – uncommitted pages (Kumar et al., 2025). In principle, EDMM allows an enclave with a 512 GB memory requirement to start with a single committed 4 KB page, although it may cause performance overhead depending on the workload (Kumar et al., 2025).
3. Motivation and Characterization
For the experiments in this section, we use a microbenchmark that performs a fixed number of ECALLs into an enclave. Each ECALL, allocates a temporary buffer on the heap, performs sequential write and read operations on the buffer, and then frees it. We measure the time to finish a fixed number of ECALL operations (system configuration shown in §7.1).
3.1. Scalability Challenges
The state-of-the-art migration mechanism for SGX (Nakashima and Kourai, 2021) is based on stop-and-copy, which has a total downtime proportional to the size of the enclave (§2.1.1). In our experiments, we observed that the total time to migrate an enclave of size (mostly heap size) 1 GB is $\approx$ 8 seconds, which increases to $\approx$160 seconds when the enclave size is 20 GB, an increase of 20$\times$. This represents a significant scaling issue, as the size of the enclave with the latest version of SGX can be scaled up to 512 GB.
A stop-and-copy method has four phases: save, transfer, initialize and restore. As shown in Figure 3, for small enclave sizes ($\approx$ 10 MB), the total migration time is dominated by the save and restore phases. However, for larger enclave sizes ($>$100 MB), the network transfer phase dominates the total migration time (up to 96% for 10 GB+ enclaves).
Insight: There is a requirement for an efficient migration scheme where the total downtime is independent of the enclave size.
Figure 3. For large enclaves, the total migration time is dominated by the enclave initialization time on the destination machine and the total transfer time.
3.2. Enclave Initialization and Down Time
After the network transfer time, the time to initialize an enclave affects the migration time the most ($17\%$, see Figure 3) since the initialization time of an enclave is proportional to the number of committed pages (§2). EDMM significantly reduces the initialization time of an enclave (and consequently the total downtime). This comes at the cost of runtime performance overheads (i.e., total migration time) (refer to (Kumar et al., 2025)).
Insight: There should be a balance between the number of committed pages, the desired initialization time and the subsequent performance overheads (see §7).
3.3. Challenges in Adopting Live Migration for SGX
We shall now briefly discuss the challenges in adopting modern live migration techniques (Clark et al., 2005; Hines et al., 2009) for Intel SGX.
3.3.1. Challenges in Pre-Copy Migration:
Tracking an application’s address space for identifying dirty pages is a key requirement for a pre-copy migration method ((§2.1). Currently, non-TEE migration methods use the dirty bit of the page table entry to track writes to a data page (cri, 2023a, b) as it is set automatically by the hardware (Doc, [n. d.]).
Prior work (Bulck et al., 2017a, b) reported that the dirty bits are set even for SGX-backed pages, and can be used to track write operations inside an enclave. However, their experiments were conducted on the Intel Skylake i7-6700 processor (Intel, 2025a), a desktop-class CPU released nine years ago. Since then, Intel has phased out SGX support on desktop CPUs, limiting SGX functionality to server-class processors (to Wikimedia projects, 2015). Furthermore, Intel has incorporated multiple security techniques to thwart attacks against Intel SGX (Randmets, 2021; Fei et al., 2021; Intel, 2025e).
We thus observed in our experiments (see Appendix A) with Xeon-6354 (Intel, 2025d), a server-class CPU that was launched in 2021, that dirty bits are not set for SGX-backed pages even after multiple write operations.
Insight: The standard page dirty-bit tracking method will not work for an SGX application, preventing a pre-copy migration.
3.3.2. Challenges in Post-Copy Migration
@@
Tracking the memory accesses of a restored application is a key requirement for a post-copy method (§2.1). A non-TEE migration process uses the userfaultfd (Doc, 2022) mechanism to capture page faults. In userfaultfd, the operating system delivers the page faults to the migration process for handling. If the accessed data is yet to be restored, the migration process fetches the data from the source machine over the network and services the fault. As of now, the userfaultfd mechanism only supports anonymous pages, huge pages, or shared memory (Doc, 2022). As the secure memory of SGX enclaves is device-backed /dev/isgx, the latest userfaultfd mechanism will not work to track memory accesses (Corbet, 2023, 2022; Doc, 2022; Kernel, 2023).
Insight: A post-copy migration based on userfaultfd will not work for SGX enclaves.
3.4. Intel SGX vs Intel TDX
Intel TDX introduces the concept of hardware-protected confidential virtual machine (CVM) (Misono et al., 2024; Kuvaiskii et al., 2024) (also known as trusted domains) that allows execution of unmodified applications within a TEE environment. TDX enables live migration of a CVM with hardware support.
Size of the TCB: The TCB (Trusted Computing Base) of TDX includes the TDX-enabled processor with Virtualization Technology (VT), Multi-key Total Memory Encryption (MKTME) and SGX (Cheng et al., 2024b). The TCB also consists of the TDX Module, which contains 20,000 lines of code (Intel, 2024), and the entire guest operating system with additional software stacks running within the CVM (Cheng et al., 2024b). Such a large TCB provides a wide attack surface and makes Intel TDX inherently susceptible to errors and vulnerabilities(Misono et al., 2024). As a matter of fact, 54% of the identified vulnerabilities related to Intel TDX are associated with the underlying firmware (Misono et al., 2024).
Initialization overhead Misono et al.(Misono et al., 2024) showed that the boot time of a TDX CVM goes up with an increase in the allocated memory – with 256 GB taking $\approx$ 24 seconds. However, we observed that a TDX CVM on an Emerald Rapid Machine (§7.1) takes $\approx$ 26 seconds, irrespective of the allocated memory – an optimization enabled by TD-shim where a background thread adds memory after CVM has booted up (Containers, 2025). However, this results in a performance degradation of the first malloc() and memset() operations (see Appendix C).
Memory encryption overhead Applications executing in a CVM have large performance overheads due to total memory encryption (Misono et al., 2024). As per Coppolino et al.(Coppolino et al., 2024), PyTorch inference executed on SGX with the LibOS Gramine (che Tsai et al., 2017) demonstrates improvements in performance of up to 78.6% compared to Intel TDX, and 32.25% compared to AMD SEV-SNP. This aligns with our experiments.
3.4.1. Advantage of SGX
SGX’s TCB consists of the processor, the microcode and firmware that provide security, and the secure part of an application’s code. The untrusted portion of the application and the host operating system are not part of the TCB. SGX protects the Ring-3 application, whereas TDX protects the entire CVM. TDX cannot protect a Ring-3 application running inside a CVM from the OS or other applications within the CVM, which is a limitation (Misono et al., 2024).
Furthermore, SGX’s flexible execution framework can address the performance concerns due to security measures. With SGX, the performance overhead in terms of memory encryption and costly enclave entry and exit are relevant only for the secure portion of the application (§2.2). The untrusted part executes in a conventional manner with no performance loss.
Entity Trusted? Access
Operating system Page fault sequence
Migration manager (CRIU) Page fault sequence
Trusted Library (ConstMigLib) Everything
Table 1. Entities involved in migration and their trust level
(a) Scope of ConstMig.
(b) A high-level overview of the working of ConstMig.

### v5

# ConstMig: Enabling Secure Live Migration of Large Intel SGX-based applications

@@ Abstract.
Cloud service providers are adopting Trusted Execution Environments, or TEEs, to provide hardware-guaranteed security to applications running on remote, untrusted data centers. However, migrating such applications still uses the decade-old stop-and-copy-based method, which introduces large downtimes. Modern live-migration approaches such as pre-copy and post-copy do not work for such applications due to limitations imposed by the TEE.
We propose ConstMig, a near-zero downtime live-migration mechanism for large memory footprint TEE-based applications. ConstMig is fully compatible with containers, virtual machines (VMs) and microVMs. Our prototype, built on Intel SGX, a TEE solution from Intel, has a near-zero downtime irrespective of enclave size. ConstMig reduces the total downtime by 77-96% for a suite of SGX applications with multi-GB memory footprints compared to state-of-the-art TEE-based migration, MigSGX.
@@ 1. Introduction
Cloud service providers, or CSPs, are adopting trusted execution environments, or TEEs, to execute security-sensitive applications (Bacon et al., 2010; Orobosade et al., 2020; Singh and Chatterjee, 2017; Baumann et al., 2014; Gribov et al., 2017; Wang et al., 2020; Priebe et al., 2018; Antonopoulos et al., 2020; Dahl et al., 2018; Hong et al., 2020; tf-encrypted, 2025; Tramer and Boneh, 2019; Shafran et al., 2021) on remote, untrusted cloud nodes. In a TEE setting, the hardware ensures the application’s security. The hypervisor and operating system (OS) are assumed to be untrusted (§2.1). Leading CSPs such as Microsoft and Google provide (Cfir Cohen, 2023; ju-shim, 2023) several TEE options such as Intel SGX (Intel, 2019c), Intel TDX (Cheng et al., 2024a), AMD SEV (Kaplan et al., 2016), ARM TrustZone (Alves and Felton, 2004) (with Confidential Compute Architecture or CCA planned (Ltd., )), and RISC-V Keystone (Lee et al., 2020).
@@
Deploying TEEs in data centers presents unique challenges in migration of applications from one machine to another. CSPs rely on efficient live migration to provide high availability, load balancing, efficient resource management, failure tolerance, seamless upgradation, and code patching (Khanna et al., 2006; Arzuaga and Kaeli, 2010; Zaw and Thein, 2012; Jin et al., 2011). In a non-TEE setting, CSPs rely on modern live migration mechanisms such as pre-copy (Clark et al., 2005) and post-copy (Hines et al., 2009) employed by hypervisor or a migration manager (§2.4) to achieve a near-zero downtime during migration (Clark et al., 2005; Hines et al., 2009). However, these techniques are not available in a TEE setting as they require access to either userspace fault handling (userfaultfd) or tracking page writes (using DIRTY bits in the page table) which are prohibited by the TEE hardware (see Section 2.4(Li and Li, 2024; Labs, ; Wang et al., 2022) as the operating system (OS), hypervisor, or any other migration manager is not a part of the trusted computing base (TCB).
Recognizing this challenge, TEE vendors integrated hardware support to enable live migrations of TEE applications. A migration manager typically orchestrates the migration and uses the API exposed by the hardware to handle the secure data pages to TEE applications – which otherwise would be inaccessible (or ineligible) to the migration manager due to TEE security guarantees. AMD enables migration of secure VMs using Migration Agents – special VMs that provide migration keys, enforce guest policies, and form part of the VM’s TCB (61). Intel TDX uses a small VM known as MigTD that assists the untrusted VMM in accessing the secure state (after encryption). ARM TrustZone does not support live migration; however, the proposed ARM CCA will have hardware-enabled live migration (Ltd., ). Intel SGX, which allows for TEE enablement at the per-application level, has no hardware support for migration. Hence, prior work in enabling migration for SGX relied on the classical stop-and-copy method for migrating SGX enclaves (Nakashima and Kourai, 2021; Liang et al., 2019; Alder et al., 2018; Guerreiro et al., 2020; Soriente et al., 2019). In the stop-and-copy method, an executing application is paused, a checkpoint is taken, and the checkpoint is transferred to the destination machine, where it is subsequently resumed. This approach has a large downtime, which is proportional to the memory footprint (see Figure 1). Given that Intel has planned support for SGX till Intel Xeon 7 – Diamond Rapids(Larabel, ), we believe that missing live migration support is a key omission which this paper aims to solve. Secure VMs provided by Intel TDX, which do have hardware support for live migration, suffer from malicious co-hosted applications, a large TCB, and a high memory encryption cost even in the unsecure portions of an application (performance overheads). SGX has a small TCB, low initialization time, and no overheads for the unsecure portion (§2.3).
@@
We present ConstMig, a novel live migration mechanism for SGX enclaves with a near-zero downtime that is independent of the memory footprint of the application. ConstMig requires no additional hardware and has far lower performance overheads than the state-of-the-art (see Figure 1). ConstMig addresses a fundamental challenge in enabling live migration for enclave-based systems: a fine-grained tracking of memory accesses on the destination machine to ensure that the restored application reads valid data from the source machine – instead of an invalid, all-zero page (§4). We propose two access tracking mechanisms: ❶ based on fine-grained page fault tracking ( §4.1) using kernel modifications and ❷ based on fine-grained pointer access tracking (§4.2). In addition to providing all the security guarantees of SGX, ConstMig also provides a freshness guarantee during the migration process, which ensures that an attacker cannot replay old encrypted enclave data during migration (§5).
Our key contributions are as follows:
@@
    We identify the key challenges faced while adopting modern migration schemes for enclaves.
@@
2. Background & Motivation
Here, we discuss some relevant background for the paper.
2.1. Intel Software Guard eXtensions or SGX
Intel Software Guard Extensions (SGX) enable secure execution of applications on untrusted machines by creating a secure sandbox called an enclave (Costan and Devadas, 2016). These enclaves are isolated from even privileged entities like the OS, hypervisor, and system administrators. At boot, SGX reserves a portion of main memory which is hardware-managed and encrypted using a hardware key inaccessible to any software component. The usable part of this reserved space is called the Enclave Page Cache or EPC (the rest is used by SGX to store metadata). The latest version of SGX, referred to as scalable SGX, supports an EPC of up to 512 GB (Intel, 2022b, a), compared to the earlier version (SGXv1) with a usable EPC of only 92 MB (Kumar et al., 2022a; Kumar and Sarangi, 2021; Intel, 2019a).
Memory management: SGX ensures an enclave’s integrity by validating the signature of a loaded enclave (Costan and Devadas, 2016) at initialization time. However, this process can become costly for enclaves with large memory footprints, as all pages must be provisioned upfront. For instance, as per our experiments, initializing a 20 GB enclave takes $\approx 30$ seconds. To address this, Intel introduced Enclave Dynamic Memory Management (EDMM) (McKeen et al., 2016), which allows developers to specify a minimal set of committed pages during initialization, while additional uncommitted pages can be added dynamically (Kumar et al., 2025). In principle, EDMM enables an enclave requiring 512 GB to start with a single committed 4 KB page, though this flexibility may introduce performance overhead depending on the workload (Kumar et al., 2025; Dhanraj et al., 2025).
extitSGX execution framework: SGX provides security at the application level. Here, an application is partitioned into secure and unsecure components. The secure component executes within an enclave, incurring performance overheads due to security guarantees, whereas the unsecure component runs conventionally without overhead (see Figure LABEL:fig:scoping). Interaction between the two occurs via enclave calls (ECALL) from the unsecure side and outside calls (OCALL) from the enclave (Costan and Devadas, 2016).
2.2. Intel Trusted Domain Extension or TDX
Intel Trusted Domain eXtension or TDX is the next-generation TEE offering from Intel which introduces the concept of a hardware-protected confidential virtual machine (CVM) (Misono et al., 2024; Kuvaiskii et al., 2024)– also known as Trusted Domain (or TD)– instead of just a single application. TDX allows unmodified applications to execute inside a hardware-protected VM (although the OS and the hypervisor must be modified to handle TDX specific interrupts (Cheng et al., 2024b; Misono et al., 2024)). TDX enables live migration of a CVM with hardware support.
Intel TDX supports dynamic memory management through the Dynamic Physical Address Metadata Table (PAMT), which stores metadata essential for securely managing physical memory assigned to Trust Domains (TDs). While the entire DRAM is encrypted using Intel Multi-Key Total Memory Encryption (MKTME), TDX enables dynamic allocation and deallocation of PAMT entries, allowing memory pages to be added to or removed from a TD at runtime, enabling full DRAM encryption and integrity protection without huge initialization overheads.
2.3. Intel SGX vs Intel TDX
Intel TDX is the latest TEE offering from Intel which aims to secure the whole VM and has hardware support for a live migration. TDX is the natural choice for deploying an application with TEE support with minimal development effort. However, SGX brings its own advantages that are either lacking or missing in TDX.
2.3.1. Execution Overhead
As discussed, Intel SGX divides an application into secure and unsecure portions. The secure portion executes in an enclave inside SGX and incurs the overhead due to security guarantees of SGX. The untrusted portion, however, suffers no additional performance overhead. However, Intel TDX does not follow such a partitioned approach. An application executes completely within a CVM and all of its data is protected by the TDX module (Misono et al., 2024).
[…]
extbfSGX Advantage: SGX’s flexible execution framework can address the performance concerns due to security measures. With SGX, the performance overhead in terms of memory encryption and costly enclave entry and exit are relevant only for the secure portion of the application (§2.1). The untrusted part executes in a conventional manner with no performance loss. Figure 2 shows the end-to-end performance of Key-Value workload for varying amount of sensitive data – for 100% everything executes in SGX and, say, for 0% only 60% data is in SGX, rest in the untrusted region outside SGX. With 100% of data as secure, SGX has a performance overhead of $\approx$12% compared to TDX. However, with 20% of data as secure, SGX outperforms TDX by $\approx$ 9% and approaches the performance of No-SGX.
2.3.2. TCB Size:
The TCB (Trusted Computing Base) of TDX includes the TDX-enabled processor with Virtualization Technology (VT), Multi-key Total Memory Encryption (MKTME) and SGX (Cheng et al., 2024b), along with TDX Module ($\approx$20 K LOC (Intel, 2024)), and the entire guest OS and software stacks within the CVM (Cheng et al., 2024b). Such a large TCB provides a wide attack surface and makes Intel TDX inherently susceptible to errors and vulnerabilities(Misono et al., 2024). Misono et al.(Misono et al., 2024) report that 54% of the identified vulnerabilities related to Intel TDX are associated with the underlying firmware. SGX’s TCB consists of the processor, the microcode and firmware that provide security, and the secure part of an application’s code. The untrusted portion of the application and the host operating system are not part of the TCB.
2.3.3. Initialization overhead
After the network transfer time, the time to initialize an enclave/CVM affects the migration time the most ($17\%$, see Figure 3). With SGX, the initialization time of an enclave is proportional to the number of committed pages – a low count of committed pages reduces the initialization time, but may incur performance overhead (Kumar et al., 2025; Dhanraj et al., 2025). TDX uses PAMT to dynamically add 4 KB pages after booting. We observed that a TDX CVM on an Emerald Rapid Machine (§6.1) takes $\approx 26$ seconds to boot up, irrespective of the allocated memory (similar observation made by Misono et al.(Misono et al., 2024)).
2.4. Migration methods and Challenges with SGX
@@
The performance of a migration method is evaluated using two metrics: total downtime – the period during which the application is inactive – and total migration time – the duration from migration start to completion (Stoyanov and Kollingbaum, 2018; Nakashima and Kourai, 2021).
Figure 3. For stop-and-copy, for large enclaves, the total migration time is dominated by the enclave initialization time on the destination machine and the total transfer time.
2.4.1. Stop-and-Copy Migration
In the stop-and-copy migration approach (Clark et al., 2005), the application is paused on the source machine while its state (e.g., registers, memory, open files, network) is transferred to the destination machine. It resumes execution only after the transfer completes. As a result, both total migration time and downtime scale are proportional to the application’s memory footprint.
[…]
Figure 3 shows the four phases of stop-and-copy migration: save, transfer, initialize and restore. For small enclave sizes ($\approx$ 10 MB), the total migration time is dominated by the save and restore phases. However, for larger enclave sizes (>>100 MB), the network transfer phase dominates the total migration time (up to 96% for 10 GB+ enclaves).
@@ 2.4.2. Pre-copy Migration
As shown in Figure LABEL:fig:pre_copy, pre-copy migration proceeds as follows: ❶ Snapshot the application’s state and track subsequent write operations. ❷ Transfer the snapshot to the destination. ❸ Identify and track modified (dirty) pages. If their count is below a threshold, ❹ pause the application and transfer the dirty pages else send the dirty pages again and go back to step three. ❺ Resume the application on the destination. Here, total migration time scales with memory footprint, while downtime depends on the number of dirty pages in the final transfer.
Challenge in adopting for SGX: Tracking an application’s address space to identify dirty pages is essential for the pre-copy migration method. Non-TEE migration approaches typically rely on the page table entry’s dirty bit, which is automatically set by the hardware upon a write to a data page (64; 66; 27). Prior work (Bulck et al., 2017b, a) has shown that dirty bits are also set for SGX-backed pages, enabling the tracking of write operations within enclaves. However, these findings were based on experiments conducted on the Intel Skylake i7-6700, a desktop-class processor released almost a decade ago (Intel, 2025a). Since then, Intel has discontinued SGX support on desktop CPUs, restricting it to server-class processors (to Wikimedia projects, 2015), and has introduced several security enhancements to mitigate SGX-related attacks (Randmets, 2021; Fei et al., 2021; Intel, 2025c).
In our experiments using the latest Xeon-based servers (Icelake and Emerald-Rapids), we observed that dirty bits are not set for SGX-backed pages, even after repeated write operations (see Appendix A for a micro-benchmark to validate this claim).
@@ 2.4.3. Post-Copy Migration
In post-copy migration (Figure LABEL:fig:post_copy), ❶ the application is first paused on the source machine, and a minimal state (CPU registers, non-pageable memory (Hines et al., 2009; Clark et al., 2005)) is sent to the destination machine. ❷ The application is then resumed on the destination, while the remaining state is transferred in the background. ❸ During the transfer, memory accesses to pages still on the source machine trigger network page faults, which are transparently handled by the OS and migration mechanism. Here, total migration time scales with memory footprint, but downtime remains constant. However, performance may degrade during migration due to costly network faults.
@@
Challenge in adopting for SGX: Tracking memory accesses of a restored application is critical for the post-copy migration. In non-TEE environments, migration frameworks use the userfaultfd mechanism (Doc, 2022) to intercept page faults. With userfaultfd, the OS forwards page faults to the migration manager, which then fetches the missing page from the source machine and services the fault. However, userfaultfd supports only anonymous memory, huge pages, and shared memory regions (Doc, 2022). As SGX enclave memory is device-backed via /dev/isgx (Corbet, 2023, 2022; Doc, 2022; Kernel, 2023), userfaultfd cannot track memory accesses within enclaves.